OpenStudy (mimijola):

Find the value of x (3x+31) + (2x-6)

OpenStudy (ivitt):

Put it in like terms : (3x+2x)+31-6

OpenStudy (mimijola):

Ok one sec

Nnesha (nnesha):

you can't find the x value without `=` sign that's just an expression (therefore you can simplify it but wouldn't be able to find `x` value )

OpenStudy (ivitt):

True.

Nnesha (nnesha):

it should be an equation in order To solve for a variable

OpenStudy (ivitt):

It it will only come up as 5x+25

OpenStudy (mimijola):

Somehow it's supposed to equal 31

OpenStudy (mimijola):

Somehow it's supposed to equal 31

Nnesha (nnesha):

do you mean like this (3x+31) + (2x-6)=31 ??

OpenStudy (mimijola):

I tried to draw it out but my tablet wont allow me to so I'm just going to close the question

Nnesha (nnesha):

can you take screenshot ?? or explain how that looks like

OpenStudy (ivitt):

get rid of 31 on both sides. 3x+2x−6=0 simplify 5x−6=0 add 6 5x=6 Then divide x= \[\frac{ 6 }{ 5 }\] 0r 1.2